OpenStudy (anonymous):

a1 = 6 an+1=-2an+3 help pleaseee!!

OpenStudy (anonymous):

I need the first 5 terms! i'm confused on how to solve this

OpenStudy (pfenn1):

Do you mean this is a series of equations where\[a _{1}=6\]\[a _{n+1}=-2a _{n}+3\]?

OpenStudy (anonymous):

yeah

OpenStudy (pfenn1):

so \[a _{1}=6\]\[a _{2}=-2a _{1}+3=-2(6)+3=-9\]\[a _{3}=?\]

OpenStudy (anonymous):

Oh... wow okay thanks can you help me with another one?

OpenStudy (pfenn1):

You should be able to do the next ones by yourself. Just use the equation....\[a _{3}=-2a _{2}+3=\]What would be the answer to this?

OpenStudy (anonymous):

21.. but i mean like okay if you're being asked for the general term, would you answer with the general term formula?

OpenStudy (pfenn1):

Yes, I think so. I'm not sure if I was asked for the "general term" I would know what they were asking about.

OpenStudy (anonymous):

Okay. for this one: 3, 12, 21, 30, 39 I was asked to write a formula for the general term and then use the formula to find the 20th term my answers were an = 3+(n-1)9 and a20 = 174.. But my teacher marked it wrong and I don't understand why. Do you mind doing it to see if you get something different?

OpenStudy (pfenn1):

I think I came up with the same answer you got. [\[a _{1}=3\]Each next term is 9 more than the term before. So the general formula would be\[a _{n}=a _{n-1}+9\] Then \[a _{20}=a _{19}+9=(a _{18}+9)+9=((a _{17}+9)+9)+9=...\]\[=a _{1}+9+9+9+9+9+9+9+9+9+9+9+9+9+9+9+9+9+9+9\]\[=a _{1}+19(9)\]which is\[a _{n+1}=a _{1}+(n-1)9\]Which is exactly what you got? So I don't know why your teacher says you are wrong. You might have to re-ask this question and see if anyone else can help.